OpenStudy (anonymous):

cos(t)=1/5 what is csc(t)

OpenStudy (jdoe0001):

\(\bf cos(t)=\cfrac{1}{5}\to \cfrac{adjacent}{hypotenuse}\to \cfrac{a=1}{c=5} \\ \quad \\ c^2=a^2+{\color{blue}{ b}}^2\implies \pm\sqrt{c^2-a^2}={\color{blue}{ b}} \\ \quad \\ recall\implies csc(\theta)=\cfrac{hypotenuse}{opposite}\to \cfrac{c}{{\color{blue}{ b}}}\)